Two trains leave Cleveland at the same time. One train travels east and the other travels west. The speed of the westbound train is 3 mph greater than the speed of the eastbound train. After 3.5 hours they are 640.5 miles apart. Find the rate of each train. Assume that the trains travel in a straight line in directly opposite directions.
Answer by JulietG(1812)

I don't blame them for trying to get out of Cleveland!
.
Let E be eastbound and W be westbound.
Here's the information you're given in the question:
3.5E + 3.5W = 640.5
W = E + 3
.
Since you know the value of W, substitute it into the first equation.
3.5E +3.5(E+3) = 640.5
Distribute
3.5E +3.5E+ 10.5 = 640.5
Add the Es
7E + 10.5 = 640.5
Subtract 10.5 from each side
7E = 630
Divide each side by 7
E = 90
.
If E = 90, then W = 93.
.
Let's put it back into the question to make certain it's correct.
In 3.5 hours, E travels 315 miles
In 3.5 hours, W travels 325.5 miles
Added together, they travel 640.5 miles
Success!